# Unsupervised Learning for Quadratic Assignment

### v1

@@ 4 Results
4.1 Data Generation
We generate synthetic instances using an Erdős-Rényi (ER) graph model following Tan and Mu (2024). To build a QAP instance of size $n$, we generate a flow matrix $\mathbf{F}\in\mathbb{R}^{n\times n}$ and the location coordinates $\mathbf{X}=(\texttt{Uniform}(0,1),\texttt{Uniform}(0,1))$. We generate uniformly random weights in $[0,1]$ for the upper triangular portion of $\mathbf{F}$, then mirror these values to create a symmetric matrix. We apply an ER graph mask with edge probability $p$ to control the sparsity of connections between facilities.
@@ 4.1 Data Generation
$\mathbf{F}_{ij}$ =$=\begin{cases}\texttt{Uniform}(0,1)&\text{if }\text{rand}()<p\text{ and }i\neq j\\ 0&\text{otherwise}\end{cases},\quad\forall i<j$ (8)
@@ 4.2 Training
$\langle\mathbb{T}\mathbf{F}\mathbb{T}^{T},\mathbf{D}\rangle,$ (10)
with the model’s hidden dimension set to either $32$ or $64$. The number of layers $l$ is set to either $2$ or $4$. For the Gumbel-Softmax operator used in Equation 5, we set $\tau\in\{1,2,3,4,5\}$ and $l\in\{10,20,30,40,50\}$. The noise scale $\gamma$ is set to either $0.001$ or $0.01$. The tanh scale is set to $\alpha=40$. We use the Adam optimizer with a learning rate of $3\times 10^{-5}$ and train for $100$ epochs. We then select the model with the best validation performance to evaluate on the test set. For problem size $100$, we set $m=100$, and for problem size $200$, we set $m=200$.
4.3 Method
@@
$0.1$ 230.086 257.877 10.7768 $2.80944\times 10^{-2}$ 958.645 1037.01 7.55682 $1.91477\times 10^{-1}$
$0.2$ 478.116 515.381 7.23057 $3.14968\times 10^{-2}$ 1968.86 2074.41 5.08819 $1.82030\times 10^{-1}$
$0.3$ 730.449 773.459 5.56073 $2.98188\times 10^{-2}$ 2987.00 3110.27 3.96332 $1.96343\times 10^{-1}$
$0.4$ 984.907 1031.92 4.55588 $3.15542\times 10^{-2}$ 4009.58 4144.96 3.26614 $1.93695\times 10^{-1}$
$0.5$ 1239.22 1288.90 3.85445 $2.83198\times 10^{-2}$ 5040.71 5184.32 2.77008 $1.92888\times 10^{-1}$
$0.6$ 1496.41 1547.62 3.30895 $2.83606\times 10^{-2}$ 6072.39 6219.82 2.37033 $1.86270\times 10^{-1}$
$0.7$ 1754.76 1805.83 2.82806 $3.20354\times 10^{-2}$ 7112.34 7257.90 2.00554 $1.91129\times 10^{-1}$
$0.8$ 2025.43 2064.49 1.89199 $2.93868\times 10^{-2}$ 8154.81 8298.41 1.73045 $1.93583\times 10^{-1}$
$0.9$ 2278.93 2320.06 1.77280 $2.78900\times 10^{-2}$ 9199.26 9336.99 1.47510 $1.89588\times 10^{-1}$
Mean 1246.48 1289.50 3.33618 $2.96619\times 10^{-2}$ 5055.96 5184.90 2.48684 $1.90778\times 10^{-1}$
@@ 4.4 Effectiveness of UL-Based Initialization
$1-\frac{\langle\mathbf{P}\mathbf{F}\mathbf{P}^{T},\mathbf{D}\rangle}{\langle\mathbf{P}_{random}\mathbf{F}\mathbf{P}^{T}_{random},\mathbf{D}\rangle},$
where $\mathbf{P}$ is the learned assignment and $\mathbf{P}_{random}$ is a random assignment. The gap percentage shows consistent improvement across all problem densities, with more significant gains observed in sparser problems. For $n=100$ with density $p=0.1$, the UL approach achieves a 10.7770% improvement over random initialization, while for $n=200$, it yields a 7.55672% improvement. As problem density increases, the gap narrows but remains positive, indicating that our approach maintains its advantage even for denser problems.
@@
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 190.129 198.931 205.433 224.427 190.154 198.888 175.675 181.297 165.349 167.250
$0.2$ 422.156 434.907 443.269 469.997 422.089 434.915 478.116 410.615 388.068 390.903
$0.3$ 663.959 679.353 688.723 720.382 663.984 679.369 640.422 650.792 623.827 627.494
$0.4$ 911.248 928.427 938.500 973.599 911.228 928.329 885.413 896.848 867.245 871.310
$0.5$ 1160.92 1179.51 1189.85 1227.30 1161.02 1179.49 1133.85 1146.22 1114.71 1119.23
$0.6$ 1415.86 1435.01 1445.24 1484.27 1415.84 1435.06 1387.92 1400.88 1368.54 1373.03
$0.7$ 1673.38 1692.78 1703.09 1742.15 1673.31 1692.70 1645.29 1658.21 1625.78 1630.24
$0.8$ 1939.07 1953.37 1972.66 2001.89 1939.01 1953.46 1910.17 1919.61 1889.13 1892.27
$0.9$ 2198.54 2213.92 2229.21 2260.26 2198.55 2213.89 2171.44 2181.57 2152.00 2155.33
Average 1175.03 1190.69 1201.78 1233.81 1175.021 1190.68 1150.26 1160.67 1132.74 1136.34
@@
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 874.188 918.224 915.225 981.426 874.146 918.342 836.399 871.458 798.478 816.504
$0.2$ 1851.52 1913.14 1908.32 1998.87 1851.68 1912.98 1799.74 1849.18 1748.06 1773.71
$0.3$ 2848.72 2921.98 2987.00 3022.16 2848.77 2921.71 2788.05 2847.12 2727.83 2758.39
$0.4$ 3856.81 3937.86 3929.77 4048.26 3856.62 3937.92 3790.02 3856.20 3724.59 3758.61
$0.5$ 4878.52 4965.70 4955.69 5082.12 4878.88 4965.68 4807.95 4879.02 4738.94 4776.09
$0.6$ 5905.06 5994.81 5984.58 6114.63 5905.33 5994.91 5832.49 5905.84 5761.67 5800.11
$0.7$ 6943.11 7032.42 7023.56 7152.33 6943.24 7032.42 6869.62 6942.84 6798.78 6836.88
$0.8$ 7988.29 8076.76 8067.27 8194.76 7988.28 8077.04 7916.56 7988.74 7846.75 7884.92
$0.9$ 9039.82 9125.34 9115.28 9237.99 9040.02 9124.94 8971.13 9041.09 8904.87 8941.46
Average 4909.56 4987.36 4979.41 5092.50 4909.66 4987.33 4845.77 4909.06 4783.33 4816.30
4.5 PLUME Search for QAP
@@
Notably, the relative improvement from using PLUME search is more significant for larger problem sizes, with $n=200$ showing consistently higher percentage improvements than $n=100$ at comparable densities. This trend holds for tabu search with different parameters. For example, with parameters $\mu=1,000$, $\kappa=25$, and $\omega=25$, the tabu search with random initialization achieves $1190.69$ on average for $n=100$, while PLUME search achieves $1175.03$, yielding a 1.31520% improvement. At $n=200$, PLUME search achieves $4909.56$ on average while random initialized tabu search achieves $4987.36$, resulting in a more substantial 1.55994% improvement. The most intensive tabu search configuration, TS$(10k,100,100)$, provides the most comprehensive exploration of the solution space and thus yields the best solution quality across all initialization methods. At this parameter setting, PLUME search further improves solution quality. For graphs with $n=100$, PLUME search achieves an average solution cost of $1132.74$ compared to $1136.34$ for random initialization, yielding a 0.316807% improvement. While this may appear modest, it is important to note that as search parameters increase, the final solutions converge closer to optimality, leaving less room for improvement. The fact that PLUME search maintains its advantage even in this setting underscores the quality of its initialization. For larger graphs with $n=200$ vertices, the benefit becomes more pronounced. PLUME search achieves an average cost of $4783.33$ compared to $4816.30$ for random initialization, yielding a 0.684550% improvement. This increasing advantage with problem size suggests that PLUME search scales well to larger problems.
Table 4: Average Time comparison of tabu search with different search parameters on QAP instances with $n=100$, in ms (×$\times 10^{-3}$ s).
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 1.55014 1.47786 2.65438 2.01335 2.02096 2.32776 3.96913 3.86332 6.89681 6.40729
$0.2$ 1.40809 1.72910 2.16947 2.11550 2.19710 2.13621 4.38935 4.10793 6.19193 6.92497
$0.3$ 2.23220 2.30824 2.32095 2.20613 2.03746 2.12630 4.06415 4.28922 6.34255 6.50180
$0.4$ 2.72354 2.91837 2.04465 2.24711 1.91479 2.45331 4.05188 4.01972 6.73373 6.90415
$0.5$ 2.57762 1.60781 2.13888 2.35198 1.97659 2.44347 4.20058 4.24511 6.87188 6.36137
$0.6$ 1.90425 1.71858 2.26572 2.21770 2.34254 1.99306 3.86804 3.87752 6.63121 6.72185
$0.7$ 1.60661 1.57086 2.11504 2.34658 2.11745 2.13709 3.66044 3.91046 6.60751 6.77780
$0.8$ 1.91665 2.06804 2.02229 2.31294 2.28833 1.97817 3.77077 3.66541 6.81585 6.88984
$0.9$ 2.42608 2.11719 2.11124 2.00790 2.31411 2.31984 3.67512 4.25899 6.28867 6.47949
Mean 2.03835 1.94623 2.20474 2.20213 2.13437 2.21280 3.96105 4.02641 6.59779 6.66317
Table 5: Average Time comparison of tabu search with different search parameters on QAP instances with $n=200$, in ms (×$\times 10^{-3}$ s).
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 2.42600 2.29224 2.26313 2.38713 2.49597 2.70196 6.85827 6.82996 11.8025 11.8930
$0.2$ 2.29030 2.33664 2.38857 2.28357 2.78697 2.75465 6.62369 6.64074 11.9798 11.8780
$0.3$ 2.34508 2.32658 2.27949 2.36495 2.65163 2.77468 6.49103 6.87804 11.9024 11.7376
$0.4$ 2.18291 2.57973 2.33208 2.16072 2.85288 2.80369 6.79153 6.76758 12.0805 11.9279
$0.5$ 2.34563 2.22272 2.54632 2.50635 2.74198 2.82831 6.59106 6.78405 12.0350 11.9095
$0.6$ 2.54630 2.29580 2.33564 2.56085 2.54337 2.97725 6.65846 6.81566 12.0519 11.6976
$0.7$ 2.27941 2.27283 2.54844 2.33566 2.48008 3.13651 7.27975 7.16379 12.1370 11.8553
$0.8$ 2.35310 2.39341 2.60996 2.48153 2.37177 2.01584 7.05032 6.80656 11.6703 11.8593
$0.9$ 2.18062 2.51180 2.41696 2.68309 2.06515 2.72407 7.16032 7.26328 11.6522 11.5438
Mean 2.32771 2.35908 2.41340 2.41821 2.55442 2.74633 6.83383 6.88330 11.92351 11.81135
Tables 4 and 5 show the execution time comparison (TS runtime) between random and UL-based initialization approaches for tabu search. The inference time for generating UL-predicted solutions is negligible: $<3\times 10^{-2}$ ms for $n=100$ and $<2\times 10^{-1}$ ms for $n=200$ (shown in Table 1). When accounting for both inference and execution time, the total computational cost of UL-based initialization remains comparable to random initialization across all parameter configurations. Overall, our results show that PLUME search effectively improves the efficiency of the tabu search algorithm, enabling the discovery of better solutions within a similar computational budget.
4.6 Compared with Other Data-driven Methods
We compare with recent work by  Tan and Mu (2024), where the authors use RL and test only on QAP instances up to size 100. We run their model on $n=100$ and $p=0.7$, using their model, we observe an average cost of $1644.37$ with an average time of $\approx 150$ ms, our PLUME TS($10010k,100,$ 100) achieves better performance with a cost of $1625.78$. Notably, our methods are substantially faster, with average time costs of just $7$ ms. It should be noted that although Tan and Mu (2024) also uses tabu search as a benchmark; they employ a Python implementation, which is not as computationally efficient as our C++ implementation of tabu search. We didn’t fine-tune the RL method and fine-tuning may yield better performance. However, the gaps are so dramatically large that even with optimization, the RL approach would remain substantially inferior.
5 Generalization
Table 6: Generalization across densities for the model trained on $n=100$, $p=0.7$ and test on $n=100$ with different densities. Random initialization shows costs using random initial assignments. UL-based initialization shows costs using assignments predicted by our neural network. TS$(\mu,\kappa,\omega)$ shows costs after running tabu search from random initialization. PLUME $\text{TS}(\mu,\kappa,\omega)$ shows costs after running tabu search from UL-predicted initialization.
Density $0.1$ $0.2$ $0.3$ $0.4$ $0.5$ $0.6$ $0.8$ $0.9$
Random Initialization 257.877 515.381 773.459 1031.92 1288.90 1547.62 2064.49 2320.06
UL-based Initialization 258.090 515.928 774.288 1032.26 1282.81 1528.17 2049.10 2320.75
$\text{TS}(1k,25,25)$ 198.931 434.907 679.353 928.427 1179.51 1435.01 1953.37 2213.92
PLUME $\text{TS}(1k,25,25)$ 198.991 434.811 679.407 928.081 1177.45 1428.55 1947.81 2213.91
$\text{TS}(10k,100,100)$ 167.250 390.903 627.494 871.310 1119.23 1373.03 1892.27 2155.33
PLUME $\text{TS}(10k,100,100)$ 167.296 390.828 627.470 871.144 1118.73 1371.62 1891.02 2155.40
Table 7: Generalization across densities for the model trained on $n=200$, $p=0.7$ and test on $n=200$ with different densities.
Density $0.1$ $0.2$ $0.3$ $0.4$ $0.5$ $0.6$ $0.8$ $0.9$
Random Initialization 1037.01 2074.41 3110.27 4144.96 5184.32 6219.82 8298.41 9336.99
UL-based Initialization 1036.78 1968.86 3110.34 4199.06 5040.71 6072.39 8284.74 9358.93
$\text{TS}(1k,25,25)$ 918.299 1913.04 2921.80 3937.95 4965.65 5994.98 8076.76 9125.18
PLUME $\text{TS}(1k,25,25)$ 918.402 1851.55 2921.83 3965.58 4878.55 5905.10 8069.67 9137.63
$\text{TS}(10k,100,100)$ 816.661 1773.70 2758.83 3758.61 4776.40 5800.17 7884.66 8941.58
PLUME $\text{TS}(10k,100,100)$ 816.572 1748.18 2758.66 3768.02 4738.89 5761.62 7882.97 8946.10
5.1 Cross-density Generalization
We further study how the model generalizes across different densities, as shown in Table 6 and 7. Using models trained on a specific edge density ($p=0.7$), our results show the best performance on nearby densities ($p=0.6$ and $p=0.8$). Model performance gradually declines as we move further from the training density, with more significant drops at the extreme values ($p=0.1$ and $p=0.9$). This suggests the model captures transferable structural patterns that work best within a reasonable proximity to its training conditions. This effect is more pronounced with TS$(1k,25,25)$ compared to configurations with more extensive search parameters. On $n=100$, with TS$(1k,25,25)$, the performance difference between PLUME and random initialization at $p=0.6$ is 0.450171%, while with TS$(10k,100,100)$, the difference narrows to 0.102693%.
Table 8: Cross-size generalization: Model trained on $n=200$ instances and tested on $n=100$ instances (both with density $p=0.7$). Random Initialization shows cost from random assignments; UL-based Initialization shows cost from UL model. TS$(\mu,\kappa,\omega)$ shows costs after running tabu search from random initialization. PLUME $\text{TS}(\mu,\kappa,\omega)$ shows costs after running tabu search from UL-predicted initialization.
Random Initialization UL-based Initialization TS $(1k,25,25)$ PLUME TS $(1k,25,25)$ TS $(10k,100,100)$ PLUME TS $(10k,100,100)$
1805.83 1795.57 1692.78 1688.43 1630.24 1629.54
Table 9: Cross-size generalization: Model trained on $n=100$ instances and tested on $n=200$ instances (both with density $p=0.7$).
Random Initialization UL-based Initialization TS $(1k,25,25)$ PLUME TS $(1k,25,25)$ TS $(10k,100,100)$ PLUME TS $(10k,100,100)$
7257.90 7233.49 7023.566 7015.50 6836.88 6830.74

### v2

@@ 4 Results
Data Generation
We generate synthetic instances using an Erdős-Rényi (ER) graph model following (Tan and Mu, 2024). To build a QAP instance of size $n$, we generate a flow matrix $\mathbf{F}\in\mathbb{R}^{n\times n}$ and the location coordinates $\mathbf{X}=(\texttt{Uniform}(0,1),\texttt{Uniform}(0,1))$. We generate uniformly random weights in $[0,1]$ for the upper triangular portion of $\mathbf{F}$, then mirror these values to create a symmetric matrix. We apply an ER graph mask with edge probability $p$ to control the sparsity of connections between facilities.
@@
$\mathbf{F}_{ij}$ $=\begin{cases}\texttt{Uniform}(0,1)&\text{if }\text{rand}()<p\text{ and }i\neq j\\ 0&\text{otherwise}\end{cases},\quad\forall i<j$ (16)
@@
$\langle\mathbb{T}\mathbf{F}\mathbb{T}^{\top},\mathbf{D}\rangle,$ (18)
with the model’s hidden dimension set to either $128$ or $256$. The number of layers $n_{\text{layers}}$ is set to $3$. For the Gumbel-Softmax operator used in Equation 13, we set $\tau=3$ and $l=100$. The noise scale $\gamma=0.01$. The tanh scale is set to $\alpha=40$. We use the AdamW optimizer with a learning rate of $3\times 10^{-5}$ and train for $300$ epochs.
Method
@@
$0.1$ 214.169 257.877 16.95 $0.5747$ 913.790 1037.01 11.88 $2.0885$
$0.2$ 457.821 515.381 11.17 $0.5686$ 1908.48 2074.41 8.00 $2.0589$
$0.3$ 709.034 773.459 8.33 $0.5672$ 2917.46 3110.27 6.20 $2.0049$
$0.4$ 960.346 1031.92 6.94 $0.5673$ 3939.86 4144.96 4.95 $1.9856$
$0.5$ 1214.74 1288.90 5.75 $0.5633$ 4963.71 5184.32 4.26 $1.9822$
$0.6$ 1472.90 1547.62 4.83 $0.5588$ 5996.06 6219.82 3.60 $1.9626$
$0.7$ 1729.11 1805.83 4.25 $0.5572$ 7035.88 7257.90 3.06 $1.9375$
$0.8$ 1989.98 2064.49 3.61 $0.5535$ 8072.96 8298.41 2.72 $1.8735$
$0.9$ 2251.18 2320.06 2.97 $0.5488$ 9122.38 9336.99 2.30 $1.9690$
Mean 1222.14 1289.50 7.20 $0.5621$ 4985.62 5184.90 5.22 $1.9736$
@@ 4.1 Effectiveness of UL-Based Initialization
$1-\frac{\langle\mathbf{P}\mathbf{F}\mathbf{P}^{\top},\mathbf{D}\rangle}{\langle\mathbf{P}_{random}\mathbf{F}\mathbf{P}^{\top}_{random},\mathbf{D}\rangle},$
where $\mathbf{P}$ is the learned assignment and $\mathbf{P}_{random}$ is a random assignment. The gap percentage shows consistent improvement across all problem densities, with more significant gains observed in sparser problems. For $n=100$ with density $p=0.1$, the UL approach achieves a 16.95% improvement over random initialization, while for $n=200$, it yields a 11.88% improvement. As problem density increases, the gap narrows but remains positive, indicating that our approach maintains its advantage even for denser problems.
@@
$n=100$ $n=200$
$\text{TS}(1k,25,25)$ $\text{TS}(10k,100,100)$ $\text{TS}(1k,25,25)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random
$0.1$ 186.255 198.931 164.509 167.250 855.216 918.224 792.157 816.504
$0.2$ 417.235 434.907 386.651 390.903 1827.18 1913.14 1740.81 1773.71
$0.3$ 659.526 679.353 623.167 627.494 2820.25 2921.98 2718.91 2758.39
$0.4$ 905.508 928.427 865.668 871.310 3830.55 3937.86 3718.45 3758.61
$0.5$ 1156.03 1179.51 1113.97 1119.23 4846.67 4965.70 4728.44 4776.09
$0.6$ 1411.77 1435.01 1368.36 1373.03 5874.34 5994.81 5752.03 5800.11
$0.7$ 1667.47 1692.78 1624.33 1630.24 6912.64 7032.42 6790.50 6836.88
$0.8$ 1928.97 1953.37 1886.45 1892.27 7950.10 8076.76 7829.94 7884.92
$0.9$ 2191.91 2213.92 2151.28 2155.33 9003.56 9125.34 8888.56 8941.46
Average 1169.41 1190.69 1131.60 1136.34 4880.06 4987.36 4773.31 4816.30
4.2 PLUME Search for QAP
We then run tabu search using different initializations. Tables 2 shows the performance of PLUME tabu search compared with tabu search with random initialization. Our experimental results demonstrate that our UL-based method effectively solves QAPs. The solutions consistently outperform random initialization across all problem sizes and density parameters. Specifically, we vary the $\texttt{evaluations}:\mu$, $\texttt{neighbourhoodSize}:\kappa$, $\texttt{maxFails}:\omega$ and test PLUME search. We show that our UL-based initialization consistently outperforms random initialization within each parameter set, and the solution quality improves with increased $\texttt{evaluations}~\mu$.
The performance gap is most pronounced in sparse problems (low $p$ values). With parameters $\mu=1{,}000$, $\kappa=25$, and $\omega=25$, PLUME search yields a 6.37% improvement at $p=0.1$, $n=100$, and a 6.86% improvement at $p=0.1$, $n=200$. This advantage diminishes as problem density increases, indicating that our neural networks are particularly effective at capturing structural patterns in sparse settings.
The relative improvement from using PLUME search is more significant for larger problem sizes, with $n=200$ showing consistently higher percentage improvements than $n=100$ at comparable densities. This trend holds for tabu search with different parameters. For example, with parameters $\mu=1,000$, $\kappa=25$, and $\omega=25$, the tabu search with random initialization achieves $1190.69$ on average for $n=100$, while PLUME search achieves $1169.41$, yielding a 1.79% improvement. At $n=200$, PLUME search achieves $4880.06$ on average while random initialized tabu search achieves $4987.36$, resulting in a more substantial 2.15% improvement. The most intensive tabu search configuration, TS(10,000,100,100)(10{,}000,100,100), provides the most comprehensive exploration of the solution space and thus yields the best solution quality across all initialization methods. At this parameter setting, PLUME search further improves solution quality. For graphs with $n=100$, PLUME search achieves an average solution cost of $1131.60$ compared to $1136.34$ for random initialization, yielding a 0.42% improvement. While this may appear modest, it is important to note that as search parameters increase, the final solutions converge closer to optimality, leaving less room for improvement. The fact that PLUME search maintains its advantage even in this setting underscores the quality of its initialization. For larger graphs with $n=200$ vertices, the benefit becomes more pronounced. PLUME search achieves an average cost of $4773.31$ compared to $4816.30$ for random initialization, yielding a 0.89% improvement. This increasing advantage with problem size suggests that PLUME search scales well to larger problems.
An important observation emerges when analyzing initialization performance at larger problem scales. For $n=200$, the UL-based initialization produces solutions of comparable quality to those obtained by a lightweight tabu search configuration, $\text{TS}(1{,}000,25,25)$. This finding highlights that the UL model does not merely provide a heuristic shortcut, but rather learns to internalize and exploit the underlying problem structure. Consequently, the UL-based initialization can be interpreted as a learned search mechanism in its own right, capable of directly yielding high-quality solutions without requiring extensive local exploration.
@@ 4.2 PLUME Search for QAP
$n=100$ $n=200$
$\text{TS}(1{,}000,25,25)$ $\text{TS}(10{,}000,100,100)$ $\text{TS}(1{,}000,25,25)$ $\text{TS}(10{,}000,100,100)$
$p$ UL random UL random UL random UL random
$0.1$ 1.55014 1.47786 6.89681 6.40729 2.42600 2.29224 11.8025 11.8930
$0.2$ 1.40809 1.72910 6.19193 6.92497 2.29030 2.33664 11.9798 11.8780
$0.3$ 2.23220 2.30824 6.34255 6.50180 2.34508 2.32658 11.9024 11.7376
$0.4$ 2.72354 2.91837 6.73373 6.90415 2.18291 2.57973 12.0805 11.9279
$0.5$ 2.57762 1.60781 6.87188 6.36137 2.34563 2.22272 12.0350 11.9095
$0.6$ 1.90425 1.71858 6.63121 6.72185 2.54630 2.29580 12.0519 11.6976
$0.7$ 1.60661 1.57086 6.60751 6.77780 2.27941 2.27283 12.1370 11.8553
$0.8$ 1.91665 2.06804 6.81585 6.88984 2.35310 2.39341 11.6703 11.8593
$0.9$ 2.42608 2.11719 6.28867 6.47949 2.18062 2.51180 11.6522 11.5438
Mean 2.03835 1.94623 6.59779 6.66317 2.32771 2.35908 11.92351 11.81135
Tables 3 show the execution time comparison (TS runtime) between random and UL-based initialization approaches for tabu search. The inference time for generating UL-predicted solutions is $0.56$ ms for $n=100$ and $1.98$ ms for $n=200$ (see Table 1). As discussed, the UL-based initialization on $n=200$ achieves competitive performance with respect to TS with parameters $\mu=1{,}000$, $\kappa=25$, and $\omega=25$ (4985.62 vs. 4987.36), while the TS runtime is $2.35$ ms. It should be noted that our UL inference is performed on GPU, whereas tabu search is executed on CPU; therefore, the runtimes are not directly comparable. Nevertheless, the shorter inference cost highlights that UL is an efficient and effective method for initialization.
4.3 Compared with Other Data-driven Methods
We compare with recent work by  (Tan and Mu, 2024), where the authors use RL and test only on QAP instances up to size 100. We run their model on $n=100$ and $p=0.7$, using their model, we observe an average cost of $1644.37$ with an average time of $\approx 150$ ms, our PLUME TS(10,000,100,10010{,}000,100,100) achieves better performance with a cost of $1624.33$. Notably, our methods are substantially faster, with average time costs within 10 ms. It should be noted that although (Tan and Mu, 2024) also uses tabu search as a benchmark; they employ a Python implementation, which is not as computationally efficient as our C++ implementation of tabu search. We didn’t fine-tune the RL method and fine-tuning may yield better performance. However, the gaps are so dramatically large that even with optimization, the RL approach would remain substantially inferior.
5 Generalization
5.1 Cross-density Generalization
We further study how the model generalizes across different densities, as shown in Table 4. Using models trained on a specific edge density ($p=0.7$), we test the model’s performance on nearby densities ($p=0.6$ and $p=0.8$). Our results suggest the model captures transferable structural patterns that work best within a reasonable proximity to its training conditions. This effect is more pronounced with TS(1,000,25,25)(1{,}000,25,25) compared to configurations with more extensive search parameters.
@@ 5.1 Cross-density Generalization
Random $\text{TS}(1{,}000,25,25)$ $\text{TS}(10{,}000,100,100)$
$n$ $p$ Random UL-based Random PLUME Random PLUME
100 0.6 1547.62 1500.26 1435.01 1420.25 1373.03 1370.25
100 0.8 2064.49 2058.36 1953.37 1950.79 1892.27 1891.46
200 0.6 6219.82 6205.04 5994.98 5984.02 5800.17 5795.99
200 0.8 8298.41 8283.00 8076.76 8060.64 7884.66 7874.34
For $n=100$ and $p=0.6$, the UL-based initialization achieves a $3.06\%$ improvement over random initialization. When applying $\text{TS}(1{,}000,25,25)$, the performance gap between PLUME and random initialization is reduced to $1.03\%$, and with the stronger configuration $\text{TS}(10{,}000,100,100)$, the gap further narrows to $0.20\%$.
[…]
Train
Size
Test
Size
Random
Initialization
UL-based
Initialization
TS
$(1k,25,25)$
PLUME TS
$(1k,25,25)$
TS
$(10k,100,100)$
PLUME TS
$(10k,100,100)$
200 100 1805.83 1754.69 1692.78 1674.92 1630.24 1625.90
100 200 7257.90 7081.93 7023.566 6934.05 6836.88 6798.40